# A Simple Framework for Open-Vocabulary Zero-Shot Segmentation

### v2

@@ 4.2 Zero-shot segmentation
CLIP-DNOiser (Wysoczańska et al., 2024) - - 32.4 34.8 62.1
LiT (Zhai et al., 2022) (ViT-B + LAION-400M) 94M 63M 29.6 38.3 48.1
LiT (Zhai et al., 2022) (ViT-B + LAION-400M) 94M 63M 31.5 39.5 51.4
@@
SimZSS (ViT-B + LAION-400M) 94M 63M 31.1 38.1 48.6
SimZSS (ViT-S + COCO Captions) 23M 40M 32.8 39.5 55.5
SimZSS (ViT-B + COCO Captions ) 94M 63M 37.2 43.5 58.4
In Table 2, we observe similar trends as in the scenario without the background class (see Tab. 1). Reported results are not as unequivocal as in the former evaluation. This is in part due to the crudeness of the background detection mechanism, which contrasts with the more sophisticated approaches used by some of the competing baselines. For instance, CLIP-DNOiser (Wysoczańska et al., 2024) relies on FOUND (Siméoni et al., 2023), whereas OVDiff (Karazija et al., 2023) uses different background prototypes for each class. Once again, training on COCO Captions leads to improved performance. When training on LAION-400M, a concept is identified in less than 15% of the samples on average. As a result, concepts rarely co-occur, making it sufficient to detect the concept without localizing it in the image to minimize loss. Therefore, the model is not trained to be less confident in the background. Exploring the possibility of enlarging the list of concepts could provide further insights.
4.3 Zero-shot classification
Table 3: Zero-shot classification on ImageNet-1k. Image-level predictions are obtained by projecting the image [CLS] token onto pre-computed text embeddings of class names. Accuracy is reported for various visual pretraining and vision-language alignment methods. $\dagger$ refers to our reproduction using DINOv2 pretrained vision backbones. The remaining results are as reported in Zhai et al. (2022).
Method Visual pretraining Backbone Pretraining dataset Alignment dataset Labels Top-1 accuracy
LiT (Zhai et al., 2022) MoCo-v3 (Chen et al., 2021) ViT-B/16 ImageNet-1k CC12M+YFCC100M 55.4
LiT (Zhai et al., 2022) DINOv1 (Caron et al., 2021) ViT-B/16 ImageNet-1k CC12M+YFCC100M 55.5
LiT (Zhai et al., 2022) AugReg (Steiner et al., 2022) ViT-B/16 ImageNet-21k CC12M+YFCC100M 55.9
LiT (Zhai et al., 2022) DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M COCO Captions 22.6
LiT (Zhai et al., 2022) DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M LAION-400M 63.6
Ours
SimZSS DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M COCO Captions 24.3
SimZSS DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M LAION-400M 64.1
@@ 4.3 Zero-shot classification
The proposed method derives from LiT (Zhai et al., 2022), noted for its excellent zero-shot classification performance on ImageNet-1k (Deng et al., 2009) among other benchmarks. Although our approach is tailored for dense downstream tasks, we investigate whether SimZSS exhibits zero-shot classification capabilities or if the introduced local consistency objective breaks this property. We follow the evaluation protocol from OpenCLIP (Ilharco et al., 2021). Class names from ImageNet-1k are contextualized with the corresponding templates from CLIP (Radford et al., 2021) and embedded through the text encoder. On the vision side, images are passed through the vision encoder, and the global representation is used to obtain class predictions via projection onto the class embeddings. The accuracy of the prediction is reported.
@@
The results reported in Table 3 reveal that SimZSS not only preserves classification capabilities comparable to LiT but also exhibits slight improvements in the low-data regime. It is worth mentioning that it takes CLIP $32$ epochs on LAION-400M to train a ViT-B/16 to reach a zero-shot accuracy of $67.0$ (see Ilharco et al. (2021), Tab. 18). Overall, while pretraining on COCO Captions excels for segmentation tasks, zero-shot classification benefits from larger datasets. This suggests that a pretraining on LAION-400M followed by finetuning on COCO Captions could optimize the benefits of both datasets.
4.4 Scalability of SimZSS
We first verify in Tables 9 and 8 that SimZSS-specific operations incur only moderate memory overhead and negligible time overhead compared to LiT. Considering this, along with the data efficiency of the proposed approach, SimZSS is orders of magnitude cheaper to train than CLIP. Secondly, we escalate the number of image-caption pairs from LAION-400M during training and report the performance on zero-shot segmentation at the end of each training phase. The results depicted in Figure 4 underscore the scalability of our method SimZSS and suggest potential gains from even larger datasets, such as LAION-5B (Schuhmann et al., 2022), or increased training epochs.
@@ 4.5 Bank of concepts
Table 4: Ablation study on the impact of concept sets. Training is conducted on COCO Captions with and without class names from Pascal VOC included in the concept set, using a ViT-B/14 vision backbone. Performance on the zero-shot segmentation task is then compared based on achieved mIoU scores. $\dagger$ refers to our reproduction using DINOv2 pretrained vision backbones.
Without background With background
Method Pascal VOC concepts Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
LiT (Zhai et al., 2022) 86.1 35.5 25.6 25.8 18.1 31.5 39.5 51.4
SimZSS 88.1 42.9 28.7 32.9 22.6 36.6 42.4 53.4
SimZSS 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
At its core, SimZSS relies on identifying noun phrases in captions and filtering them through a bank of concepts. This filtering ensures that the identified noun phrases are valid candidates for visual concepts and can be used to query the images. A potential question that arises is whether the segmentation improvements of the proposed method are limited to the classes included in the bank of concepts or if they generalize to any class. To investigate this, we train SimZSS with the class names from Pascal VOC (Everingham et al., 2012) removed from the bank of concepts and report the resulting zero-shot segmentation mIoU scores across all downstream datasets.
Most importantly,  Table 4 shows that SimZSS outperforms LiT on Pascal VOC both with and without background. This invalidates the hypothesis that only classes contained in the bank of concepts benefit from the proposed local consistency objective. Furthermore, the performance on other downstream datasets only slightly deviates from the one obtained using the complete list of concepts, which further supports our claim, considering that the labels from Pascal VOC account for approximately $22\%$, $26\%$, and $33\%$ of the labels from COCO Object (Caesar et al., 2018), Cityscapes (Cordts et al., 2016), and Pascal Context (Mottaghi et al., 2014), respectively. Overall, these results suggest that our method effectively trains the text encoder to align with local features.
Alternatively,  Figure 1 suggests that the vision tower, here a ViT-B/14 pretrained with DINOv2 (Oquab et al., 2023), could handle more concepts than those present in the concept bank. Indeed, some visual concepts, such as bike (third row, columns 1-5), appear well-segmented in the image but are not highlighted in the caption, meaning the concept is absent from the concept bank, and thus, no constraint is enforced. This claim is further supported by the fine-grained vision-language alignment depicted in Figure 3. The visualization further shows that, even though concepts such as forest, chef, water buffalo, or bike are not in the concept bank, the text concepts and the corresponding patches appear well aligned.
@@ 5 Conclusion
We introduce SimZSS, a simple framework to endow pretrained pure vision models with open-vocabulary segmentation capabilities. This approach is versatile: it can accommodate various backbones, pretraining methods, and datasets, irrespective of their scale and degree of curation. Overall, SimZSS is both a computationally and data-efficient method that yields state-of-the-art results across standard zero-shot segmentation benchmarks.
Aknowledgement.
This research is supported by the Personalized Health and Related Technologies (PHRT), grant number 2021/344. Additional funding is provided by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101021347). This work is supported by a grant from the Swiss National Supercomputing Centre (CSCS) on the Swiss share of the LUMI system under project ID 606.
@@
Appendix A Experimental setup
A.1 Pretraining
A.1.1 Pretraining datasets
A.1.2 Networks architectures
A.1.3 Optimization
For COCO Captions, we conduct training over 10 epochs using a global batchsize of 16,384. We incorporate a warm-up strategy spanning 10% of the training steps, linearly ramping up the learning rate until it reaches its peak value, chosen from the set {8e-5, 3e-5, 8e-6, 3e-6}111scaled by batchsize / 256. Subsequently, we employ a cosine decay schedule for the remaining steps. Similarly, for LAION-400M, we train for 1 epoch with a global batchsize of 32,768, and we set the learning rate from the options {3e-5, 8e-6, 3e-6, 8e-7, 3e-7}. The remaining optimization settings align with those of OpenCLIP (Ilharco et al., 2021).
A.1.4 Text concepts
Appendix B Additional experiments
B.1 Impact of the vision tower
@@
In Tables 5 and 6, we explore the performance of various pretrained vision towers in zero-shot segmentation tasks. Our observations indicate that SimZSS is versatile and can integrate diverse models and pretraining approaches effectively. As expected, models with superior performance on pure vision tasks also show enhanced results when used with our method, as well as with LiT. Most importantly, regardless of the vision backbone employed, SimZSS consistently surpasses LiT. Additionally, employing denoised ViTs (Yang et al., 2024), which exhibit better semantic feature correlation, leads to improved segmentation. This outcome is anticipated since our approach assumes that features corresponding to semantically related regions are highly correlated (see Sec. 3.2).
B.2 SimZSS-specific hyperparameters
@@
Table 7: Ablation study on the impact of the loss weight $\mathbf{\lambda}$ and the temperature parameter $\mathbf{\tau}$. Training is performed on COCO Captions using a ViT-B/14 vision backbone. The performance is evaluated through a comparison of mIoU scores on the zero-shot segmentation task.
@@
Thanks to its simplicity, SimZSS has few hyperparameters, with the primary ones being $\tau$, which regulates the temperature used to query the dense visual representation, and $\lambda$, which modulates the contribution of the dense loss to the overall objective (see Eqs. 10 and 4). After a grid search on $\lambda$, $\tau$, and the learning rate, we find that the best-performing setting for training on COCO Captions is $(\lambda=0.05,\tau=0.1)$. In Table 7, we report the resulting zero-shot segmentation performance when varying $\lambda$ and $\tau$ around this optimal point.
@@
B.3 High-level profiling & runtime comparison

### v3

@@ 4.3 Zero-shot segmentation
CLIP-DINOiser (Wysoczańska et al., 2023) - - 32.4 34.8 62.1
LAION-400M
CLIP (Radford et al., 2021) (ViT-B) 0 157M 6.9 5.2 11.1
LiT (Zhai et al., 2022) (ViT-B) 94M 63M 29.6 38.3 48.1
SimZSS (ViT-B) 94M 63M 31.1 38.1 48.6
COCO Captions
LiT (Zhai et al., 2022) (ViT-B) 94M 63M 31.5 39.5 51.4
SimZSS (ViT-S) 23M 40M 32.8 39.5 55.5
SimZSS (ViT-B) 94M 63M 37.2 43.5 58.4
In Table˜2, we observe similar trends as in the scenario without the background class (see Tab.˜1). Reported results are not as unequivocal as in the former evaluation. This is in part due to the crudeness of the background detection mechanism, which contrasts with the more sophisticated approaches used by some of the competing baselines. For instance, CLIP-DINOiser (Wysoczańska et al., 2023) relies on FOUND (Siméoni et al., 2023), whereas OVDiff (Karazija et al., 2023) uses different background prototypes for each class. Once again, training on COCO Captions leads to improved performance. When training on LAION-400M, a concept from the bank of concepts is identified in less than 15% of the samples on average. As a result, concepts rarely co-occur in a single sample, making it sufficient to detect the concept without localizing it in the image to minimize the loss. Therefore, the model is not trained to be less confident in the background.
4.4 Zero-shot classification
Table 3: Zero-shot classification. Image-level predictions are obtained by projecting the image [CLS] token onto pre-computed text embeddings of class names. Accuracy is reported for various visual pretraining and vision-language alignment methods. $\dagger$ refers to our reproduction using DINOv2 pretrained vision backbones. The remaining results are as reported in Zhai et al. (2022).
Method Visual pretraining Backbone Pretraining dataset Alignment dataset Alignement samples Labels ImageNet-1k Average
LiT (Zhai et al., 2022) MoCo-v3 (Chen et al., 2021) ViT-B/16 ImageNet-1k CC12M+YFCC100M - 55.4 -
LiT (Zhai et al., 2022) DINOv1 (Caron et al., 2021) ViT-B/16 ImageNet-1k CC12M+YFCC100M - 55.5 -
LiT (Zhai et al., 2022) AugReg (Steiner et al., 2022) ViT-B/16 ImageNet-21k CC12M+YFCC100M - 55.9 -
LiT (Zhai et al., 2022) DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M COCO Captions 4M 22.6 24.4
LiT (Zhai et al., 2022) DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M LAION-400M 400M 63.6 37.5
CLIP (Radford et al., 2021) - ViT-B/16 - LAION-400M 12.8 B 67.0 47.2
@@ 4.4 Zero-shot classification
SimZSS DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M COCO Captions 4M 24.3 26.1
SimZSS DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M LAION-400M 400M 64.1 38.9
SimZSS DINOv2 (Oquab et al., 2023) ViT-B/14 LVD-142M LAION-400M 1.6B 69.3 41.3
Table 4: Training without the bank of concepts. The impact of the concept bank is evaluated by training SimZSS on COCO Captions without it, that is, the concept-level objective is enforced for all noun phrases in the captions. The training is conducted with a ViT-B/14 pretrained with DINOv2. $\dagger$ indicates results reproduced using DINOv2 pretrained vision backbones.
Without background With background
Method Bank Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
LiT (Zhai et al., 2022) - 86.1 35.5 25.6 25.8 18.1 31.5 39.5 51.4
SimZSS 89.2 41.1 27.6 31.2 20.0 35.2 42.6 56.8
SimZSS 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
@@
The proposed method derives from LiT (Zhai et al., 2022), noted for its excellent zero-shot classification performance on ImageNet-1k (Deng et al., 2009) among other benchmarks. Although our approach is tailored for dense downstream tasks, we investigate whether SimZSS exhibits zero-shot classification capabilities or if the introduced local consistency objective breaks this property. We follow the evaluation protocol from OpenCLIP (Ilharco et al., 2021). Class names are contextualized with the corresponding templates from CLIP (Radford et al., 2021) and embedded through the text encoder. On the vision side, images are passed through the vision encoder, and the global representation is used to obtain class predictions via projection onto the class embeddings. We evaluate on 38 datasets and report the average top-1 accuracy in Table˜3. The accuracy on ImageNet-1k is also reported specifically. The complete zero-shot classification results on the 38 datasets are depicted in Figure˜5.
@@
The results reported in Table˜3 reveal that SimZSS not only preserves classification capabilities comparable to LiT but also exhibits slight improvements in the low-data regime, namely when training on COCO Captions. It is worth mentioning that the performance of CLIP after 1 epoch is approximately $43\%$ (as depicted here). Overall, while pretraining on COCO Captions excels for segmentation tasks, zero-shot classification benefits from larger datasets. This suggests that a pretraining on LAION-400M followed by finetuning on COCO Captions could optimize the benefits of both datasets.
@@ 4.5 Bank of concepts
At its core, SimZSS relies on identifying noun phrases within captions and filtering them through a bank of concepts. This filtering ensures that the identified noun phrases are valid candidates for visual concepts and can be used to query the images. A key question that arises is whether the segmentation improvements of the proposed method are confined to the classes within the concept bank or if they generalize to other classes as well. To test this hypothesis, we trained SimZSS on COCO Captions without the concept bank.
As expected, the scores reported in Table˜4 are slightly lower compared to using the concept bank. However, the results consistently outperform the baseline LiT, which focuses solely on enforcing consistency between caption and image representations. Moreover, we observe that the performance of SimZSS without the concept bank is closer to the version with the concept bank than to LiT. This demonstrates that SimZSS’s performance is driven more by the consistency objective between noun phrases and localized image regions than by the concept bank itself. Considering the results achieved by concurrent baselines on the same tasks (see Tabs.˜2 and 1), we conclude that SimZSS yields great performance compared to existing methods, irrespective of whether a concept bank is used, which shows that a concept does not need to appear in the concept bank to be properly localized at test time.
Alternatively, Figure˜1 suggests that the vision tower, here a ViT-B/14 pretrained with DINOv2 (Oquab et al., 2023), could handle more concepts than those present in the concept bank. Indeed, some visual concepts, such as bike (third row, columns 1-5), appear well-segmented in the image but are not highlighted in the caption, meaning the concept is absent from the concept bank, and thus, no constraint is enforced. This claim is further supported by the fine-grained vision-language alignment depicted in Figure˜3. The visualization further shows that, even though concepts such as forest, chef, water buffalo, or bike are not in the concept bank, the text concepts and the corresponding patches appear well aligned.
@@ 5 Conclusion
We introduce SimZSS, a simple framework to endow pretrained pure vision models with open-vocabulary segmentation capabilities. This approach is versatile: it can accommodate various backbone sizes, pretraining methods, and datasets, irrespective of their scale and degree of curation. Overall, SimZSS is both a computationally and data-efficient method that yields excellent results across standard zero-shot segmentation benchmarks while preserving efficiency at inference time.
Reproducibility Statement
We will release our code and pretrained models to the public upon acceptance. The release will include a README file with detailed instructions for setting up the environment and running the training. All datasets used in our work are publicly available.
Acknowledgement
This project is funded by the Personalized Health and Related Technologies (PHRT), grant number 2021/344, as well as the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101021347). This work is also partially supported by funding from the Flemish Government under the “Onderzoeksprogramma Artificiele Intelligentie (AI) Vlaanderen”. This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) on the Swiss share of the LUMI system under project ID 606.
@@ Acknowledgement
Appendix A Additional experiments
A.1 Impact of the vision tower
@@
In Tables˜5 and 6, we explore the performance of various pretrained vision towers in zero-shot segmentation tasks. Our observations indicate that SimZSS is versatile and can integrate diverse models and pretraining approaches effectively. As expected, models with superior performance on pure vision tasks also show enhanced results when used with our method, as well as with LiT. Most importantly, regardless of the vision backbone employed, SimZSS consistently surpasses LiT. Additionally, employing denoised ViTs (Yang et al., 2024), which exhibit better semantic feature correlation, leads to improved segmentation. This outcome is anticipated since our approach assumes that features corresponding to semantically related regions are highly correlated (see Sec.˜3.2).
As shown in Tables˜5 and 6, using a ViT-L/14 as the vision backbone does not improve overall performance compared to the ViT-B/14, whether for LiT (Zhai et al., 2022) or SimZSS. This outcome cannot be attributed to overfitting due to the increased number of trainable parameters in the text encoder. Replacing the text encoder with the architecture tailored for the ViT-B/14 did not result in any performance gains. The results for the ViT-L/14 in Tables˜5 and 6 were obtained using average pooling of the dense visual representation. This approach was necessary because we observed that the [CLS] token in this model was poorly aligned with the remaining tokens, making the training of both LiTand SimZSS unsuccessful. These findings align with the lower results reported by Pariza et al. (2024a) when using frozen features from DINOv2’s ViT-L/14 for dense downstream tasks, as well as with the growing body of literature on the denoising of DINOv2 models (Yang et al., 2024; Wang et al., 2024).
A.2 Comparison with training-free approaches
We compare SimZSS with training-free approaches on the zero-shot segmentation benchmark. The results in Table˜7 show that these methods, particularly ProxyCLIP (Lan et al., 2024), can achieve strong performance without requiring any training. Moreover, these methods retain CLIP’s zero-shot classification capabilities.
Table 7: Comparison with training-free approaches. Training-free approaches are compared to SimZSS trained on COCO Captions with different vision backbones pre-trained with DINOv2. Results for the training-free approaches are reported as in Lan et al. (2024).
Without background With background
Method Backbone Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
CLIPSurgery (Li et al., 2023c) ViT-Base - - 21.9 31.4 - 29.3 - -
GEM (Bousselham et al., 2024) ViT-Base - - - - 15.7 32.6 - 46.2
MaskCLIP (Zhou et al., 2022a) ViT-Base 74.9 26.4 16.4 12.6 9.8 23.6 20.6 38.8
SCLIP (Wang et al., 2025) ViT-Base 80.4 34.2 22.4 32.2 16.1 30.4 30.5 59.1
ProxyCLIP (Lan et al., 2024) ViT-Base 80.3 39.1 26.5 38.1 20.2 35.3 37.5 61.3
MaskCLIP (Zhou et al., 2022a) ViT-Large 29.4 12.4 8.8 11.5 7.2 11.7 7.2 23.3
SCLIP (Wang et al., 2025) ViT-Large 69.1 22.3 17.6 18.6 10.9 22.3 25.0 43.5
ProxyCLIP (Lan et al., 2024) ViT-Large 83.1 37.7 25.6 40.1 22.6 34.5 39.2 60.6
SimZSS ViT-Small 87.2 37.3 23.8 29.2 17.9 32.8 39.5 55.5
SimZSS ViT-Base 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
SimZSS ViT-Large 83.9 36.4 24.5 31.1 18.8 31.9 34.6 46.6
While SimZSS does require training, the cost is manageable as it only takes 15 minutes on 8 GPUs (for training on COCO Captions) and offers several advantages. SimZSS enables the use of smaller models, such as a ViT-S/14, can be tailored to specific domains, and is generally less computationally intensive at inference time, which can effectively amortize the training cost.
A.3 Identifying bottlenecks
Table 8: Few-shot vs. zero-shot segmentation comparison. We compare the zero-shot segmentation performance of SimZSS with an alternative approach that constructs a bank of visual concepts from the training set rather than relying on embedded class names (i.e., textual concepts). For reference, we also include the performance of linear segmentation and dense nearest-neighbor retrieval, as reported in Pariza et al. (2024b).
Without background With background
Method Background Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
Dense k-NN Concept - - - - 38.7 - - 76.2
Linear Concept - - 57.6 - 40.3 - 78.3 79.8
Few-shots Concept 91.4 51.0 31.3 49.7 31.7 43.7 27.7 51.6
Few-shots Confidence 91.4 51.0 31.3 49.7 31.7 43.6 42.4 62.6
SimZSS Confidence 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
To pinpoint potential areas for improvement, we take a close look at the performance of SimZSS in the context of zero-shot segmentation, directly comparing it to other segmentation methods on the same datasets and backbone. To dissect the influence of vision-language alignment on overall performance, we strip out the language model’s class name embeddings and replace them with visual embeddings. These visual embeddings, derived from the training set, are computed by average pooling of the corresponding local representations for each class.
In Table˜8, we report the results of the above-described experiment using the ViT-B/14 from DINOv2. For datasets that include a background class, we explore two strategies: i) treating the background as a distinct class with its own concept, and ii) classifying a given patch or pixel as background when the confidence of the prediction across all foreground classes is low.
The results support the hypothesis that the background is, in general, too heterogeneous to be well captured by a single concept, even in the absence of a modality gap. We observe that the few-shots strategy generally outperforms the zero-shot one. The high variability of the difference in mIoU scores between two strategies across different datasets indicates that the modality gap is not the sole reason for the observed discrepancies. In the context of zero-shot segmentation, the concepts are common to all datasets, while in few-shot segmentation, they are dataset-specific. This is well-reflected by the similarity of the two strategies on COCO-Stuff and COCO-Object, which can be attributed to the alignment of the pretraining dataset (COCO Captions) with the downstream datasets. Overall, this suggests that the image-caption pairs dataset is a stronger limiting factor than the modality gap itself. Leveraging image-captioning methods (Li et al., 2023a; Liu et al., 2024) to generate captions for target datasets could prove beneficial. The performance achieved through dense nearest-neighbor retrieval or linear segmentation using the same vision backbone on the same datasets indicates that there is still plenty of room for improvement.
Conversely, Table 10 from DINOv2 (Oquab et al., 2023) demonstrates significantly superior performance on the linear segmentation downstream evaluation across all their models, including ViT-S/14, compared to OpenCLIP ViT-G/14 (Ilharco et al., 2021). This underscores the limitations of the cross-modal contrastive objective in learning effective visual representations. Moreover, it suggests that methods relying on pretrained CLIP-based models face a lower performance ceiling compared to our approach. Indeed, zero-shot segmentation can be viewed as pixel- or patch-level classification using a frozen linear layer initialized with the embeddings of textual labels.
A.4 SimZSS-specific hyperparameters
@@
Table 9: Ablation study on the impact of the loss weight $\bm{\lambda}$ and the temperature parameter $\bm{\tau}$. Training is performed on COCO Captions using a ViT-B/14 vision backbone. The performance is evaluated through a comparison of mIoU scores on the zero-shot segmentation task.
@@
Thanks to its simplicity, SimZSS has few hyperparameters, with the primary ones being $\tau$, which regulates the temperature used to query the dense visual representation, and $\lambda$, which modulates the contribution of the dense loss to the overall objective (see Eqs.˜10 and 4). After a grid search on $\lambda$, $\tau$, and the learning rate, we find that the best-performing setting for training on COCO Captions is $(\lambda=0.05,\tau=0.1)$. In Table˜9, we report the resulting zero-shot segmentation performance when varying $\lambda$ and $\tau$ around this optimal point.
@@
A.5 Scalability of SimZSS
We first verify in Tables˜13 and 12 that SimZSS-specific operations incur only moderate memory overhead and negligible time overhead compared to LiT. Considering this, along with the data efficiency of the proposed approach, SimZSS is orders of magnitude cheaper to train than CLIP. Secondly, we escalate the number of image-caption pairs from LAION-400M during training and report the performance on zero-shot segmentation at the end of each training phase. The results depicted in Figure˜4 underscore the scalability of our method SimZSS and suggest potential gains from even larger datasets, such as LAION-5B (Schuhmann et al., 2022), or increased training epochs.
A.6 Image resolution
Table 10: Ablation study on the impact of the training resolution. Training is conducted on COCO Captions with an image resolution of either $224\times 224$ or $448\times 448$ pixels using a ViT-B/14 vision backbone pretrained with DINOv2. Performance on the zero-shot segmentation task is then compared based on achieved mIoU scores. $\dagger$ refers to our reproduction using DINOv2 pretrained vision backbones.
Without background With background
Method Resolution Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
LiT (Zhai et al., 2022) $448\times 448$ 84.2 33.9 25.5 24.9 17.3 31.1 39.7 50.5
SimZSS $448\times 448$ 87.2 40.2 27.9 29.4 20.4 35.2 43.2 56.6
LiT (Zhai et al., 2022) $224\times 224$ 86.1 35.5 25.6 25.8 18.1 31.5 39.5 51.4
SimZSS $224\times 224$ 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
Segmentation methods typically benefit from higher image resolutions during training. To assess whether this holds for SimZSS, we conducted training on COCO Captions, increasing the image resolution from $224\times 224$ pixels to higher resolution images, $448\times 448$ pixels. For that experiment, the vision backbone utilized was a ViT-B/14 pretrained with DINOv2. Apart from the increased resolution, all other training procedures remain unchanged.
As shown in Table˜10, neither LiT nor SimZSS exhibited any performance gains when trained with higher-resolution images. This can likely be attributed to the fact that higher resolutions reduce the receptive field of each patch, leading to poorer semantic alignment between local representations and the global image representation, typically captured by the [CLS] token. In LiT Zhai et al. (2022), this misalignment hampers zero-shot segmentation performance, as it weakens the alignment between local visual features and the global textual representation, which is fundamental to the task. The same trend is observed with SimZSS, which builds upon the knowledge acquired from the alignment of global textual and visual representations.
We further explore the effect of using higher resolutions during evaluation. After training with an input resolution of $224\times 224$, we assess zero-shot segmentation performance using a target size for the shorter side of either 448 or 896. As shown in Table˜11, increasing the evaluation resolution proves beneficial primarily for the Cityscapes (Cordts et al., 2016) dataset. This is expected, given that Cityscapes consists of high-resolution images that undergo substantial downscaling, leading to a greater potential loss of information.
Table 11: Ablation study on the impact of the evaluation resolution. Training is conducted on COCO Captions with an image resolution of $224\times 224$ pixels using a ViT-B/14 vision backbone pretrained with DINOv2. Zero-shot segmentation performance is reported for images resized to a shorter side of either 448 or 896 pixels. $\dagger$ refers to our reproduction using DINOv2 pretrained vision backbones.
Without background With background
Method Shorter side Pascal VOC Pascal Context COCO-Stuff Cityscapes ADE20K Pascal Context COCO-Object Pascal VOC
LiT (Zhai et al., 2022) $448$ 86.1 35.5 25.6 25.8 18.1 31.5 39.5 51.4
SimZSS $448$ 90.3 43.1 29.0 33.0 21.8 37.2 43.5 58.4
LiT (Zhai et al., 2022) $896$ 79.7 34.0 23.8 30.4 20.6 31.0 37.2 49.2
SimZSS $896$ 86.1 43.3 28.9 36.1 22.7 37.7 41.7 59.9
A.7 High-level profiling & runtime comparison